OpenStudy (anonymous):

log _x 9 = -2 x is the base i get the question.. its just when you have x^-2 = 9... it becomes 1/x^2 = 9... then HOW does it become x^2 = 1/9 ? what is done there?

OpenStudy (anonymous):

cross multiply: 9* x^2 = 1 divide both sides by 9: x^2 = 1/9

OpenStudy (anonymous):

oh i see thank you!!

OpenStudy (anonymous):

\[1 /x ^{2} = 9\]

OpenStudy (anonymous):

yw